# Motion-R1: Chain-of-Thought Reasoning and Reinforcement Learning for Human Motion Generation

### v3

@@ 4.4 Qualitative Analysis and Visualization
Out-of-Distribution Visualization. To evaluate generalization beyond the training distribution, we present qualitative comparisons under two types of out-of-distribution captions, as shown in Figure  4 (middle and right).
The first prompt describes a complex, multi-stage action with implicit causal logic: “After hearing a loud noise, a person turned around, stepped back cautiously with hands raised defensively, and then slowly approached.” Motion-R1 captures this chain of reactions in a temporally coherent manner, clearly separating reaction, retreat, and re-approach. In contrast, MotionLLM either merges steps or lacks defensive gestures, resulting in an incomplete or disordered motion sequence.
The second prompt—“A person is serving in badminton”—involves abstract semantics and task-specific dynamics. Motion-R1 synthesizes a plausible serve-like motion, including arm lifting, body shift, and a forward striking gesture, resembling real badminton actions. MotionLLM, however, produces repetitive or generic motions lacking physical intent or domain alignment.
These results demonstrate that Motion-R1 generalizes well to unseen instructions by leveraging explicit CoT-style reasoning and planning, while baselines such as MotionLLM struggle with long-horizon dependencies or abstract concepts.
4.5 Ablation Study
To assess the contribution of each component in our framework, we conduct an ablation study on the HumanML3D dataset, as shown in Table 2. Specifically, we evaluate the impact of three key components: CoT, the semantic similarity reward ($R_{\text{sem}}$), and the motion similarity reward ($R_{\text{motion}}$).
All ablated variants achieve comparable performance in most metrics, reflecting the robustness of our design. Notably, enabling CoT alone results in the poorest performance, with significantly lower R-Precision scores (Top-1: $0.340{\pm}.002$) and higher FID ($0.530{\pm}.015$) and MM-Dist ($4.216{\pm}.015$), indicating insufficient alignment between generated motions and the input descriptions. Adding either $R_{\text{sem}}$ or $R_{\text{motion}}$ to CoT yields a substantial boost, particularly in R-Precision Top-1 (increasing to $0.483{\pm}.002$) and FID (dropping to $0.281{\pm}.008$), suggesting that semantic and motion-level feedback are both crucial for enhancing generation quality.
[…]
5 Conclusion and Limitations
Conclusion   In this work, we presented Motion-R1, a unified framework for text-to-motion generation that combines Chain-of-Thought reasoning and reinforcement learning. Our approach decomposes complex language into structured action plans and optimizes motion synthesis using GRPO, improving controllability, diversity, and generalization. To support reasoning-based training, we introduced the MotionCoT Data Engine for automatic CoT-style annotation.
Limitations   While Motion-R1 shows promising results, it still has limitations. The CoT decomposition is generated by prompting general-purpose LLMs, which may introduce noise or suboptimal planning in ambiguous instructions. Additionally, although GRPO simplifies the RL pipeline by removing the need for explicit value functions or preference data, it still requires careful design of multi-component rewards. Future work will explore adaptive reward learning and interactive feedback to further enhance generation quality.
Appendix A MORE QUALITATIVE VISUALIZATIONS
@@
To further highlight the strengths of Motion-R1, we provide additional qualitative examples that emphasize its performance in capturing semantic alignment, producing realistic motions, and handling out-of-distribution (OOD) prompts. These visualizations complement the main results by demonstrating the model’s ability to interpret complex or abstract instructions, maintain temporal consistency, and reflect user intent. All results are generated without post-processing. More examples and detailed CoT traces are provided in the supplementary video.
Appendix B Details of Prompt
@@
B.1 MotionCoT Data Engine
This prompt type is used during the construction of the MotionCoT dataset. It encourages the language model to generate structured reasoning traces and explicit motion planning steps based on free-form motion descriptions.
@@ B.2 Motion-R1
B.3 Motion-R1 w/o CoT
This form is used for training w/o CoT. No reasoning steps are involved, and the model directly maps textual descriptions to motion token sequences. This prompt type is structurally simple and helps the model learn basic text-to-motion alignment without reasoning supervision.
@@ B.3 Motion-R1 w/o CoT
Appendix C Experimental Settings
C.1 Datasets
We conduct experiments on widely-used dataset for text-driven motion generation, HumanML3D [3]. HumanML3D [3] is the most extensive dataset currently available for 3D human motion conditioned on natural language. It comprises $14,616$ motion samples and $44,970$ corresponding textual descriptions, featuring a vocabulary of 5,371 unique words. The motion data originates from AMASS [52] and HumanAct12 [53], and is uniformly processed. All sequences are adjusted to 20 frames per second, cropped to a maximum duration of 10 seconds, and retargeted to a canonical skeleton facing the positive Z-axis. Each motion is annotated with at least three distinct textual descriptions, with an average description length of 12 words.
C.2 Evaluation Metrics
Following [3, 46, 9], our evaluation metrics are summarized as follows:
@@ C.2 Evaluation Metrics
3) Diversity. To assess the variation within generated motion sequences, we follow the protocol introduced by Guo et al. [3]. Specifically, we randomly sample $S_{\text{dis}}$ pairs of motions from the generated set, and compute the average $\ell_{2}$ distance between each pair’s motion features. Let $f_{\text{pred},i}$ and $f^{\prime}_{\text{pred},i}$ denote the feature embeddings of the $i$-th sampled pair, the diversity is defined as:
@@ C.3 Training Details
Appendix D Experiments Compute Resources
@@
Appendix E Broader impacts
This work proposes Motion-R1, a reasoning-guided and reinforcement-optimized framework for text-to-motion generation. It has the potential to benefit applications in animation, virtual reality, embodied AI, and assistive technologies.
As with other generative models, there is a risk of misuse, such as synthesizing inappropriate or misleading motion sequences. We discourage such use and emphasize that Motion-R1 is intended solely for research and beneficial applications.

### v4

@@ 4.4 Ablation Study
5 Conclusion
In this work, we introduced Motion-R1, a unified framework for text-to-motion generation that integrates a Decomposed CoT Data Engine and RL Binding. The Decomposed CoT Data Engine generates structured, step-by-step reasoning traces to decompose complex language into interpretable action plans, while RL Binding optimizes motion synthesis through GRPO with semantic and motion-level reward alignment. Together, these innovations enable coherent, semantically faithful, and high-quality motion generation without relying on costly human annotations, improving controllability, diversity, and generalization.
6 Ethics Statement
This work introduces Motion-R1, a framework for text-to-motion generation. While the method enables high-fidelity and semantically aligned motion synthesis, it shares general risks associated with generative models, including the potential misuse for creating misleading, unsafe, or inappropriate motion sequences. We strongly discourage such applications and emphasize that Motion-R1 is intended strictly for research, educational, and other socially beneficial purposes. Users should exercise caution, ensure compliance with ethical guidelines, and consider privacy and consent when applying the model to real-world scenarios.
7 Reproducibility Statement
We are committed to ensuring the reproducibility of our work. Details of the framework, training protocols, and hyperparameters are provided in Section 3, Subsection 4.1 and Appendix A.3;  A.4. Code for our method will be made publicly available to support replication and future research.
Appendix A Appendix
A.1 Statement on the Use of LLM
Large Language Models (ChatGPT by OpenAI) were used exclusively to improve the clarity and fluency of English writing. They were not involved in research ideation, experimental design, data analysis, or interpretation. The authors take full responsibility for all content.
A.2 More Qualitative Visualizations
@@
To further highlight the strengths of Motion-R1, we provide additional qualitative examples that emphasize its performance in capturing semantic alignment, producing realistic motions, and handling out-of-distribution (OOD) prompts. As shown in Figure 4, Motion-R1 successfully interprets complex or abstract instructions, maintains temporal consistency, and reflects user intent. All results are generated without post-processing. More examples are provided in the supplementary video.
A.3 Details of Prompt
@@
A.3.1 Decomposed CoT Data Engine
The following prompt is used in the Decomposed CoT Data Engine, which encourages the language model to generate structured reasoning traces and explicit motion planning steps based on free-form motion descriptions.
@@ A.3.2 Motion-R1
A.3.3 Motion-R1 w/o decomposed CoT
This form is used for training w/o decomposed CoT. No reasoning steps are involved, and the model directly maps textual descriptions to motion token sequences. This prompt is structurally simple and helps the model learn basic text-to-motion alignment without reasoning supervision.
@@ A.3.3 Motion-R1 w/o decomposed CoT
A.4 Experimental Settings
A.4.1 Evaluation Metrics
Following Guo et al. (2022a); Zhang et al. (2023a); Wu et al. (2024), our evaluation metrics are summarized as follows:
@@ A.4.1 Evaluation Metrics
3) Diversity. To assess the variation within generated motion sequences, we follow the protocol introduced by Guo et al. Guo et al. (2022a). Specifically, we randomly sample $S_{\text{dis}}$ pairs of motions from the generated set, and compute the average $\ell_{2}$ distance between each pair’s motion features. Let $f_{\text{pred},i}$ and $f^{\prime}_{\text{pred},i}$ denote the feature embeddings of the $i$-th sampled pair, the diversity is defined as:
@@ A.4.2 Training Details
A.4.3 Experiments Compute Resources
@@
A.5 Qualitative Results on BABEL
Results on BABEL. Table 3 reports results on the BABEL Punnakkal et al. (2021) dataset, which contains long, multi-label activity annotations. Motion-R1 outperforms prior methods across all metrics. It achieves the highest R-Precision ($\mathbf{0.536}{\pm}.004$), surpassing InfiniDreamer ($\underline{0.522}{\pm}.008$) and other baselines, indicating stronger text–motion semantic alignment. In terms of motion fidelity, Motion-R1 attains a substantially lower FID of $\mathbf{0.53}{\pm}.006$, nearly halving the best baseline ($\underline{1.14}{\pm}.05$ from DoubleTake/InfiniDreamer). For motion–language embedding distance, Motion-R1 also sets the best score ($\mathbf{6.16}{\pm}.141$), suggesting better cross-modal consistency. Moreover, its Diversity ($\mathbf{8.90}{\pm}.095$) exceeds even ground-truth motion ($8.52$), reflecting the ability to synthesize varied yet realistic motions.
Overall, these results demonstrate that Motion-R1 generalizes effectively to BABEL, producing semantically accurate, diverse, and high-fidelity motions under complex multi-label activity settings.
Table 3: Quantitative results of Motion-R1 on BABEL Punnakkal et al. (2021). The evaluations are conducted 20 times to obtain a $95\%$ confidence interval. Best results are highlighted in bold and the second best in underline.
Methods R-Precision $\uparrow$ FID $\downarrow$ $M$-Dist $\downarrow$ Diversity $\uparrow$
Ground Truth $629^{\pm.001} 0.0$$0004^{\pm.00}$ $3.51^{\pm.01}$ $8.52^{\pm.09}$
TEACH Athanasiou et al. (2022) $0.461^{\pm.012}$ $1.43^{\pm.04}$ $93^{\pm.01} 7.7$$71^{\pm.11}$
DoubleTake Shafir et al. (2023) $0.483^{\pm.009}$ $\underline{1.14}^{\pm.05}$ $6.97^{\pm.01}$ $\underline{8.28}^{\pm.09}$
DiffCollage Zhang et al. (2023b) $0.487^{\pm.009}$ $1.83^{\pm.05}$ $6.74^{\pm.01}$ $7.89^{\pm.11}$
InfiniDreamer Zhuo et al. (2024) $\underline{0.522}^{\pm.008}$ $\underline{1.14}^{\pm.11}$ $\underline{6.35}^{\pm.01}$ $7.97^{\pm.05}$
Motion-R1 $\textbf{0.536}^{\pm.004}$ $\textbf{0.53}^{\pm.006}$ $\textbf{6.16}^{\pm.141}$ $\textbf{8.90}^{\pm.095}$
A.6 Limitations
Despite its advantages, Motion-R1 has limitations. The Decomposed CoT Data Engine relies on general-purpose LLMs, which may produce noisy or suboptimal plans under ambiguous instructions. Furthermore, while RL Binding streamlines policy optimization, careful design of motion- and semantic-level rewards remains crucial. Future work will explore adaptive reward learning and interactive feedback mechanisms to further enhance motion quality and robustness.